HUBBLE SPACE TELESCOPE DAILY REPORT #3096 PERIOD COVERED: 0000Z (UTC) 04/16/02 - 0000Z (UTC) 04/17/02 Daily Status Report as of 107/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.2 Completed Five Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.3 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.4 Completed FGS/1 9234 (Calibrating the Mass-Luminosity Relation at the End of the Main Sequence) Fine Guidance Sensor #1 was used to calibrate the mass-luminosity relation for stars less massive than 0.2 Msun, with special emphasis on objects near the stellar/brown dwarf border. The proposal completed with no reported errors. 1.5 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.6 Completed ACS/WFC 9020 (Preliminary ACS Sensitivity) The Advanced Camera for Surveys (HFC) was used to observe a spectrophotometric standard star through each filter of each camera to assess the sensitivity of the instrument. The star is placed at the center of the aperture, and two images are taken through each filter. There were no reported problems. 1.7 Completed Two Sets of STIS/MA1/MA2 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.8 Completed STIS/CCD/MA2 9051 (Identifying Damped Lyman-alpha Galaxies at z~1) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to look for damped Lyman-alpha absorption systems that contain the bulk of the neutral gas in the Universe in the redshift range z = 0.5 - 5, yet the nature of the galaxies responsible for the absorption is not well understood. Only recently have observers found more than a handful of damped absorbers at redshifts z < 1.5. Using the FIRST Bright Quasar Survey {FBQS}, with over a 1000 quasars, the proposers have undertaken a survey to build a complete picture of he nature of the galaxies responsible for damped Lyman-alpha absorption systems at z~1 and to double the sample size at this redshift. No problems were encountered. 1.9 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 1.10 Completed STIS/CCD 9110 (A Search for Kuiper Belt Object Satellites) The Space Telescope Imaging Spectrograph (CCD) was used to investigate whether the large number of collisions thought to have taken place in the primordial Kuiper belt suggest that many Kuiper belt objects {KBOs} could have suffered binary-forming collisions similar to that which formed the Pluto -- Charon binary. Detection of such KBO satellites would allow measurement of KBO masses, would help to understand the past collisional environment of the Kuiper belt, and would give a context to the otherwise unique-seeming formation of the Pluto -- Charon binary. The proposal completed with no reported problems. 1.11 Completed STIS/CCD/MA2 8919 (MAMA Sensitivity and Focus Monitor C10) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to monitor the sensitivity of each MAMA grating mode to detect any change due to contamination or other causes, and also to monitor the STIS focus in a spectroscopic and an imaging mode. All observations were successful. 1.12 Completed STIS/CCD/MA1 9184 (A Survey for Missing Baryons in Highly Ionized Intergalactic Gas at Low Redshift) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to observe six additional low-z QSOs with the STIS FUV E140M echelle mode {7 km s^-1 resolution}. Combined with archival data, this will increase the sample redshift path by a factor of ~7 compared to the published data. With the echelle data, we will {1} measure the number of O VI absorbers per unit redshift {dN/dz} and their minimum cosmological mass density with a limiting equivalent width of W_Lambda ~50 mAngstrom , {2} examine whether the O VI absorption arises in photoionized, collisionally ionized, or multiphase gas, and {3} study the dependence of the O VI system properties on environment. In addition to testing this prediction regarding the location of the missing baryons, the data will have applications to many other topics such as low-z LyAlpha absorbers and the physical properties and abundances of gas in the Milky Way halo. There was no reported anomalous behavior. 1.13 Completed STIS/MA1 8921 (MAMA Fold Distribution) The Space Telescope Imaging Spectrograph (MA1) was used to monitor the performance of MAMA microchannel plates using a MAMA fold analysis procedure. The fold analysis provides a measurement of the distribution of charge cloud sizes incident upon the anode giving some measure of changes in the pulse-height distribution of the MCP and, therefore, MCP gain. All observations completed normally. ************************************************** The following scheduled proposal did not execute due to the BALL decision to delay the ACS high voltage on commanding: ACS/SBC 9008 (MAMA Initial Turn-on and Anomalous Recovery Procedure) ************************************************** The following proposals were not successful due to the Gyro #3 event when all subsequent acquisitions/re-acquisitions failed: WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) Two Sets of STIS/CCD 9131 (Imaging the Host Galaxies of High Redshift Type Ia Supernovae) ************************************************** 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 2 Scheduled Re-acquisitions: 8 Successful: 3 Per HSTAR 8610, the acquisition at 106/162546Z failed to RGA control. All subsequent acquisitions/re-acquisitions also failed. The gyro #3 motor current event (HSTAR 8609) may have contributed to the failures (see 2.3). 2.2 FHST Updates: Scheduled: 14 Successful: 13 The roll delay update at 107/071956Z failed when the error box results showed a value for QDVERRV0 of 356.130. This is related to the Gyro #3 event. HSTAR 8611 was written.. 2.3 Operations Notes: Using ROP DR-18A, the 486 engineering status buffer limits were adjusted twelve times. Per HSTAR 8609, the Gyro #3 motor current (GRG2_3MC) flagged out-of-limits high at 106/181052Z with a value of 204.4. The parameter returned to normal bounds at 106/101058Z. Following the Gyro #3 event and the subsequent acquisition/re-acquisition failures, the following ROPs were executed: ROP DF-14A for RGA bias uplinks at 106/2108Z. 107/0046Z, 107/233Z, and 107/0719Z ROP IC-2 to turn on SSA-1 for the interval 107/0107Z to 107/0153Z; to turn on SSA-2 for the intervals 107/050831Z to 107/052955Z, 107/075709Z to 107/081611Z, and 107/085941Z to 107/092257Z. ROP IC-1 to turn on MATX-1 at 107/041316Z (no acquisition of signal) and for the interval 107/041437Z to 107/041936Z; to turn on MATX-2 for the interval 107/082018Z to 107/092257Z. For the Gyro #3 event, the following were performed according to operations requests: Gyro Bias updates at 106/2108Z, 107/0046Z, 107/0233Z and 107/0719Z. Attitude reference updates at 106/2148Z, 107/0416Z and 107/0905Z. During the execution of the final attitude reference update at 107/0905Z, per HSTAR 8612, a 486 status buffer message was received indicating that the system momentum limit had been exceeded. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Gyro #1 will be powered on this morning, and the gyro configuration will be modified to reflect active gyros 3, 4, 2, 1, with gyro #3 in the shadow mode. This is expected to lead to the resumption of acquisitions/re-acquisitions and consequently science operations. A Battery #3 capacity test is scheduled to begin at approximately 107/1124Z. Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW